OpenStudy (anonymous):

FACTOR COMPLETLY: 36X^2-1= 21X^2-26X+8= 8X^3-49X^2-49x= 5X^2-34x-7= X^2-X-132= 2X^3-14X^2+24X= X^2+11XY+24Y^2= X^2+X-72

OpenStudy (anonymous):

can someone please help me with this?

OpenStudy (anonymous):

36x^2-1=(6x-1)(6x+1) 21x^2-26x+8=(7x-4)(3x-2) 8x^3-49x^2-49x=x(8x^2-49x-49)=x(8x+7)(x-7)

OpenStudy (anonymous):

5x^2-34x-7=(5x+1)(x-7) x^2-x-132=(x-12)(x+11)

OpenStudy (anonymous):

2x^3-14x^2+24x=2x(x^2-7x+12)=2x(x-3)(x-4) x^2+11xy+24y^2=(x+8y)(x+3y)

OpenStudy (anonymous):

x^2+x-72=(x+9)(x-8)

OpenStudy (anonymous):

last 2, i rally appreciate your help!! 49x^2+25= 256m^4-81=

OpenStudy (anonymous):

is it +25 or -25?

OpenStudy (anonymous):

256m^4-81=(16m^2-9)(16m^2+9)-(4m-3)(4m+3)(16m^2+9)

OpenStudy (anonymous):

+25

OpenStudy (anonymous):

can u deal with complex numbers?

OpenStudy (anonymous):

i have the optio to choose that is prime also

OpenStudy (anonymous):

the answer for 256m^4-8 is all that after the = sign?

OpenStudy (anonymous):

sorry instead of the minus there must be an equal symbol

OpenStudy (anonymous):

so is just this =(4m-3)(4m+3)(16m^2+9)

OpenStudy (anonymous):

the only way i know how to factorize 49x^2+25 is: (7x-5i)(7x+5i)

OpenStudy (anonymous):

kk perfect i will try that way (:

OpenStudy (anonymous):

good luck..and if u find any mistake,you may keep it:)

OpenStudy (anonymous):

i just cheked all the answers they are correct, 98% :p i missed 1 because of a symbol! than you a LOT!